OpenStudy (anonymous):

Which equation represents the line passing through the points (2, −2) and (−5, 12)? 2x − y = 2 2x + y = 2 x + 2y = 2 x − 2y = 2

OpenStudy (mertsj):

Can you find the slope ?

OpenStudy (mertsj):

\[m=\frac{-2-12}{2-(-5)}=\frac{-14}{7}=-2\]

OpenStudy (mertsj):

\[y-(-2)=-2(x-2), y+2=-2x+4, 2x+y=2\]

OpenStudy (mertsj):

The second one.